# Competitive Equilibrium for Chores: from Dual Eisenberg-Gale to a Fast, Greedy, LP-based Algorithm

### v1

@@ 3. The Convex Maximization Program
This is surprising because past work (Bogomolnaia et al., 2017; Boodaghians et al., 2022) has repeatedly stressed that a challenge for computing CE for chores is that iterative methods that work in the interior are attracted to non-CE points. Our new program completely circumvents this issue. The remainder of the paper will be concerned with developing a provably fast and highly practical iterative method for computing (approximate) KKT points of Chores Dual Redundant. We note, though, that Chores Dual Redundant opens the door to many other interior optimization methods, and we hope, for example, that efficient interior-point methods can be developed for it, as well as new first-order methods for large-scale problems.
4. Greedy Frank-Wolfe Method
@@
In this section we introduce a simple first-order optimization method for computing chores CE based on Chores Dual Redundant. The method only requires solving an LP at every iteration, and we will see that it has many desirable properties. As above, let $Y$ be the convex polytope of solutions to Chores Dual Redundant, and let $y=(\beta,p)\in Y$ denote elements of $Y$. Let $f$ be the Chores Dual Redundant objective (after dropping $\sum_{j}p_{j}$, which sums to the constant $\sum_{i}B_{i}$), i.e.
$f(y)=-\sum_{i}B_{i}\log\beta_{i}$
The algorithm we will use is a greedy variant of the Frank-Wolfe method (FW) (Frank and Wolfe, 1956). The greedy Frank-Wolfe (GFW) method generates a sequence of iterates $\{y_{t}\}_{t}$, each contained in $Y$. Each iterate is generated by linearizing the objective, and greedily maximizing the linerization over $Y$ to generate the next iterate. This yields the following algorithm,
(6) $y^{t+1}=\arg\max_{y\in Y}\langle\nabla f(y^{t}),y-y^{t}\rangle$
@@
For convex minimization, the FW method has been immensely popular, for example in machine learning application (Jaggi, 2013), and it has previously been used to compute CE in various settings with goods (Gao and Kroer, 2020; Panageas et al., 2021). In the goods setting, one must be careful about stepsize selection, in order to guarantee convergence; typically a step is taken from the current iterate in the direction of the point maximizing (or minimizing in the convex case) the linearization. Interestingly, in the chores setting this is not the case; because we are maximizing a convex objective, we can greedily move directly to the point returned by the linear maximization. GFW has previously been studied for generic convex maximization problems in the context of certain machine learning applications (Mangasarian, 1996; Journée et al., 2010). It is instructive to briefly understand why we can be greedy with this method. Because we have a convex function, the following inequality holds,
(7) $f(y^{t+1})-f(y^{t})\geq\langle\nabla f(y^{t}),y^{t+1}-y^{t}\rangle.$
@@
Note that the right-hand side is exactly the objective in Eq. 6, and thus, since $y^{t}$ is a feasible solution, the RHS is weakly greater than zero. Moreover, the solution strictly improves, unless we have already found a local maximum (i.e. a KKT point and thus a CE). This immediately implies that we have a monotonically-improving sequence of iterates. Secondly, note that there always exists a vertex of $Y$ which is optimal, and since there is a finite set of vertices in $Y$, this implies that the method converges in finite time. These observations were formalized by Mangasarian (1996):
@@
Let $f$ be a differentiable convex function that is bounded above on a polyhedron $Y$. The GFW algorithm generates a sequence of iterates $\{y^{t}\}_{t\geq 1}$ such that
@@
    The algorithm terminates with a $y^{T}$ satisfying first order optimality conditions in a finite number of iterations $T$.
While the above guarantees finite-time convergence, it does not give a rate at which this convergence occurs. Later, some convergence-rate results were shown by Journée et al. (2010). Most related to our chores CE setting, they show that for strongly convex functions, GFW has a best-iterate convergence rate of $O(1/\sigma_{f}\epsilon^{2})$, where the convergence is in terms of some iterate $k\in\{1,\ldots,T\}$ having a small stepsize $\|y^{k+1}-y^{k}\|\leq\epsilon$, and $\sigma_{f}$ is the strong convexity parameter. However, our objective is not strongly convex, and thus we cannot apply the results from Journée et al. (2010) to the chores CE setting. In the next sections, we will show how to generalize the result of Journée et al. (2010) to a class of Bregman-strongly-convex functions. We then show that our objective indeed satisfies this generalized notion of strong convexity, and thus is guaranteed a small multiplicative stepsize difference. We go on to show that this implies convergence to chores CE. In the appendix, we also give an alternative proof of this convergence directly from arguing about CE structure, without going through Bregman strong convexity.
4.1. General Convergence Results for $\log$-Type Convex Maximization
In this part, we generalize our Chores EG dual problem and show general convergence results for $\log$-type convex maximization over any polyhedral set. We show that the multiplicative difference between two consecutive iterates is guaranteed to be smaller than any $\epsilon>0$ after $O(\frac{1}{\epsilon^{2}})$ number of iterations. We then show in the following section that such multiplicative “stepsize convergence” implies convergence to an approximate chores CE.
@@ 4.1. General Convergence Results for $\log$-Type Convex Maximization
(9) $0\leq\langle\nabla f(x^{t}),x^{t+1}-x^{t}\rangle\leq f(x^{t+1})-f(x^{t})\leq f^{*}-f(x^{t})<+\infty,$
@@
(10) $\langle\nabla f(x^{t}),x^{t+1}-x^{t}\rangle\geq 0\quad\forall t=0,1,\ldots,T.$
@@
(11) $f(x^{t+1})\geq f(x^{t})+\langle\nabla f(x^{t}),x^{t+1}-x^{t}\rangle+\alpha D_{h}(x^{t+1}\,\|\,x^{t})\quad\forall t=0,1,\ldots,T.$
Summing Eq. 11 up over $t=0,\ldots,T$ and applying Eq. 10, we have
(12) $\alpha T\min_{0\leq t\leq T}D_{h}(x^{t+1}\,\|\,x^{t})\leq\alpha\sum_{t=0}^{T}D_{h}(x^{t+1}\,\|\,x^{t})\leq f(x^{T+1})-f(x^{0})\leq f^{*}-f(x^{0}).$
@@
(13) $D_{IS}(y\,\|\,x)=D_{h}(y\,\|\,x)=\sum_{i}\left(-\log{\frac{y_{i}}{x_{i}}}+\frac{y_{i}}{x_{i}}-1\right),$
where $D_{IS}(\cdot,\|\,\cdot)$ refers to the Itakura–Saito distance. Then, for the $\log$-type function in Eq. 8, we trivially have
(14) $g(y)\geq g(x)+\langle\nabla g(x),y-x\rangle+\min_{i}a_{i}\;D_{IS}(y\|x),$
@@
(15) $\frac{1}{3}\sum_{i}\left(\frac{y_{i}}{x_{i}}-1\right)^{2}\leq D_{IS}(y\,\|\,x)\quad\text{if }-\log{\frac{y_{i}}{x_{i}}}+\frac{y_{i}}{x_{i}}-1\leq c:=\sqrt{3}-1-\frac{1}{2}\log{3}\quad\forall\;i.$
@@
4.2. Convergence to Approximate Competitive Equilibrium
@@
[…]
Lemma 5 ().
@@
($t$-LP) $\min_{p,\beta\geq 0}\quad$ $\sum_{i}B_{i}\frac{\beta_{i}}{\beta^{t-1}_{i}}$
@@
Let $x$ and $\mu$ denote the dual variables corresponding to the first and second constraints of Eq. $t$-LP, respectively. We use $(p^{t},\beta^{t})$ and $(x^{t},\mu^{t})$ to denote the optimal primal and dual variables at the $t$-th iteration, respectively.
By KKT conditions and strong duality of LP, $(p,\beta)$ is an optimal solution to Eq. $t$-LP if and only if there exists a pair of $(x,\mu)$ such that $(p,\beta,x,\mu)$ satisfies
@@
Recall that $(p^{t},\beta^{t})$ is an optimal solution to Eq. $t$-LP. By condition (ii), we have $\frac{p^{t}_{j}}{d_{ij}}=\beta^{t}_{i}\geq\frac{p^{t}_{j^{\prime}}}{d_{ij^{\prime}}}$ for all $j^{\prime}\in[m]$ if $x^{t}_{ij}>0$ for all $i,j$. This guarantees every buyer is allocated her optimal bundle subject to the amount of budget she earns.
@@
(18) $\sum_{i}x^{t}_{ij}=-\mu=\frac{\sum_{i}B_{i}\frac{\beta_{i}}{\beta^{t-1}_{i}}}{\sum_{i}B_{i}}\quad\text{if }p^{t}_{j}>0.$
By 4, this holds for all $j\in[m]$. It follows that
(19) $\sum_{i}\bar{x}^{t}_{i}=\frac{\sum_{i}B_{i}}{\sum_{i}B_{i}\frac{\beta^{t}_{i}}{\beta^{t-1}_{i}}}\sum_{i}x^{t}_{i}=1\quad\mbox{for all }j\in[m].$
@@
Therefore, the lemma follows by the definition of $\epsilon$-strongly approximate equilibrium. ∎
Therefore, by Lemmas 5 and 3, we can conclude that the GFW algorithm finds an $\epsilon$-approximate CEEI in $\left\lceil\frac{3(g^{*}-g(x^{0}))}{\min_{i}a_{i}}\frac{1}{\epsilon^{2}}+\frac{g^{*}-g(x^{0})}{c\min_{i}a_{i}}\right\rceil$ iterations. Observe that $g^{*}-g(x^{0})$ can be upper bounded by $\sup_{p,\beta}-\sum_{i}B_{i}\log(\beta_{i})$. To this end, note that for each agent $i$, $1/\beta_{i}\leq\min_{i}d_{ij}/p_{j}$ (implied by the constraints of our program). Since $\sum_{j}p_{j}=\sum_{i}B_{i}$, this implies that there exists at least one chore $j$, such that $p_{j}\geq\sum_{i}B_{i}/m$, implying that $d_{ij}/p_{j}\leq mD/(\sum_{i}(B_{i}))$, where $D=\max_{i,j}d_{ij}$. Therefore, we have
$-\sum_{i}B_{i}\log(\beta_{i})$ $=\sum_{i}B_{i}\log(1/\beta_{i})$
$\leq\sum_{i}B_{i}\log(mD/(\sum_{i}B_{i}))$ $\leq n\max_{i}B_{i}\log(mD/\min_{i}B_{i})$
Theorem 6 ().
@@
The FW algorithm finds an $\epsilon$-approximate CEEI in $\left\lceil\frac{3(g^{*}-g(x^{0}))}{\min_{i}a_{i}}\frac{1}{\epsilon^{2}}+\frac{g^{*}-g(x^{0})}{c\min_{i}a_{i}}\right\rceil\in\mathcal{O}(n\max_{i}B_{i}\log(mD/\min_{i}B_{i})/\epsilon^{2})$ iterations.
@@
For CEEI, our results (Theorem 6) improve the current state-of-the-art convergence rate in terms of the number of iterations relative to $\epsilon$: The EPM by Boodaghians et al. (2022) requires $\mathcal{\tilde{O}}(n^{3}/\epsilon^{2})$ iterations, and the combinatorial algorithm by Chaudhury et al. (2022) requires $\mathcal{\tilde{O}}(nm/\epsilon^{2})$ iterations, in contrast to $\mathcal{{\tilde{O}}}(n/\epsilon^{2})$ iterations of GFW [^10] For CEEI, note that $\max_{i}B_{i}=\min_{i}B_{i}=1$., to reach an approximate CE. Furthermore, we expect the cost of each iteration to be cheaper for GFW. In GFW, we only need to solve an LP in every iteration. Moreover, the approximate solution of the LP poses no numerical issues. In contrast, solving the convex QP in EPM is a more challenging task. It is hard to find exact running times for general convex QPs, and even quite good approximate solutions can lead to serious numerical issues, as we show in Section 5. Further, we believe that the LP solved in every iteration of GFW can be solved (in worst case) using algorithms faster than the ones used for general-purpose LPs (van den Brand, 2020), and we leave this as an interesting question for future research.
@@
The result in Theorem 6 can also be recovered with mildly stronger guarantees through a primal-dual analysis of the GFW (See Appendix A). The analysis in Appendix A gives us further insights on each run of GFW: for instance, in each iteration of GFW where we do not yet have a good approximate CE point, we make a large improvement on the objective function (Claim 4).
5. Experiments

### v2

@@ 3 The Chores Dual Program
This is surprising because past work (Bogomolnaia et al. 2017, Boodaghians et al. 2022) has repeatedly stressed that a challenge for computing CE for chores is that iterative methods that work in the interior are attracted to non-CE points. Our new program completely circumvents this issue. The remainder of the paper will be concerned with developing a provably fast and highly practical iterative method for computing (approximate) KKT points of (Chores Dual Redundant). We note, though, that (Chores Dual Redundant) opens the door to many other optimization methods, and we hope, for example, that efficient interior-point methods can be developed for it, as well as new first-order methods for large-scale problems.
4 Greedy Frank-Wolfe Method
@@
In this section we introduce a simple first-order optimization method for computing chores CE based on (Chores Dual Redundant). The method only requires solving an LP at every iteration, and we will see that it has many desirable properties.
As above, let $Y$ be the convex polyhedron of the feasible set of (Chores Dual Redundant), and let $y=(\beta,p)\in Y$ denote elements of $Y$. Let $f$ be the (Chores Dual Redundant) objective (after dropping $\sum_{j}p_{j}$, which sums to the constant $\sum_{i}B_{i}$), i.e.,
$f(y)=-\sum_{i}B_{i}\log\beta_{i}.$
The algorithm we will use is a greedy variant of the Frank-Wolfe (FW) algorithm (Frank and Wolfe 1956). For a general differentiable minimization problem over a closed convex set, say $\min_{x\in\mathcal{X}}\tilde{f}(x)$, the FW method conducts the following steps iteratively. In every iteration $t$, the FW algorithm finds a minimizer $s_{t}$ to the linear approximation of the function $\tilde{f}$ around the previous iterate point $x_{t-1}$. Thereafter, the algorithm optimally chooses a step $\tau_{t}$ towards $s_{t}$ from $x_{t-1}$ and sets $x_{t}=x_{t-1}+\tau_{t}(s_{t}-x_{t-1})$ for $\tau_{t}\in[0,1]$. The advantage of this method over other first-order methods like gradient descent is that it avoids projection (which can often be expensive).
1:An initial feasible point in the dual space $y^{0}=(\beta^{0},p^{0})\in\mathbb{R}^{n+m}_{\geq 0}$ such that $\sum_{j}p^{0}_{j}=\sum_{i}B_{i}$
2:for $t=1,2,\ldots$ do
3:     $y^{t+1}=\arg\max_{y\in Y}\,\langle\nabla f(y^{t}),y-y^{t}\rangle$
4:     if $\langle\nabla f(y^{t}),y^{t+1}-y^{t}\rangle=0$ then return $y^{t}$
5:     end if
6:end for
Algorithm 1 Greedy Frank-Wolfe (GFW) algorithm
@@
For convex minimization, the FW method has been immensely popular, for example in machine learning application (Jaggi 2013), and it has previously been used to compute CE in various settings with goods (Gao and Kroer 2020, Panageas et al. 2021). In the goods setting, one must be careful about stepsize selection, in order to guarantee convergence; typically a step is taken from the current iterate in the direction of the point maximizing (or minimizing in the convex case) the linearization. Interestingly, in the chores setting this is not the case; because we are maximizing a convex objective, we can greedily move directly to the point returned by the linear maximization. In particular, for our (Chores Dual Redundant) program, we can greedily choose our current iterate point to be the maximizer $s_{t}$ of the linear approximation of the function $f$ around $y_{t-1}$. That is, we set $y_{t}=s_{t}$ where $s_{t}={\operatorname{\mathrm{arg\,max}}}_{s\in Y}\langle\nabla f(y_{t-1}),s-y_{t-1}\rangle$. This yields 6. We call this algorithm the Greedy Frank Wolfe (GFW) algorithm.
GFW has previously been studied for generic problems of maximizing a convex function in the context of certain machine learning applications (Mangasarian 1996, Journée et al. 2010). It is instructive to briefly understand why we can be greedy with this method. Because we have a convex function, the following inequality holds,
$f(y^{t+1})-f(y^{t})\geq\langle\nabla f(y^{t}),y^{t+1}-y^{t}\rangle.$ (9)
@@
Note that the right-hand side is exactly the objective in the line 2 of 6, and thus, since $y^{t}$ is a feasible solution, the RHS is weakly greater than zero. Moreover, the solution strictly improves, unless we have already found a local maximum (thus a KKT point and a CE). This immediately implies that we have a monotonically-improving sequence of iterates. Secondly, note that there always exists a vertex of $Y$ which is optimal, and since there is a finite set of vertices in $Y$, this implies that the method converges in finite time. These observations were formalized by Mangasarian (1996):
@@
Let $f$ be a differentiable convex function that is bounded above on a polyhedron $Y$. Then, 6 generates a sequence of iterates $\{y^{t}\}_{t\geq 1}$ such that
@@
    The algorithm terminates with a $y^{T}$ satisfying the first-order optimality condition
    $\langle\nabla f(y^{T}),y-y^{T}\rangle\leq 0\quad\text{for all }y\in Y$
    in a finite number of iterations $T$.
Figure 2: The trajectory of GFW until termination on a $2$-agent-$8$-chore instance, starting from a feasible interior point. The gray region is the feasible set projected onto the $(\beta_{1},\beta_{2})$ subspace. The blue lines with arrows denote the trajectory of GFW. The orange arrows and dotted lines denote gradient ascending directions of $f$ and corresponding planes perpendicular to these directions, respectively.
In Fig. 2, we illustrate a trajectory of the GFW algorithm on an instance with 2 agents and 8 chores. The earning requirements of agents are set to be one and disutilities are drawn from the [0, 1] uniform distribution.
While the above guarantees finite-time convergence, it does not give a rate at which this convergence occurs. Later, some convergence-rate results were shown by Journée et al. (2010). Most related to our chores CE setting, they show that for strongly convex functions, GFW has a best-iterate convergence rate of $O(1/(\sigma_{f}\epsilon^{2}))$, where the convergence is in terms of some iterate $k\in\{1,\ldots,T\}$ having a small stepsize $\|y^{k+1}-y^{k}\|\leq\epsilon$, and $\sigma_{f}$ is the strong convexity parameter. However, our objective is not strongly convex, and thus we cannot apply the results from Journée et al. (2010) to the chores CE setting. In the next sections, we will show how to generalize the result of Journée et al. (2010) to a class of Bregman-strongly-convex functions. We then show that our objective indeed satisfies this generalized notion of strong convexity, and thus is guaranteed a small multiplicative stepsize difference. We go on to show that this implies convergence to chores CE. In the appendix, we also give an alternative proof of this convergence directly from arguing about CE structure, without going through Bregman strong convexity.
4.1 General GFW Convergence Results for Convex Maximization
In this part, we generalize our Chores EG dual problem and show general convergence results for $\log$-type convex maximization over any nonempty closed convex set. We show that the multiplicative difference between two consecutive iterates is guaranteed to be smaller than any $\epsilon>0$ after $O(\frac{1}{\epsilon^{2}})$ number of iterations. We then show in the following section that such multiplicative “stepsize convergence” implies convergence to an approximate chores CE.
@@ 4.1 General GFW Convergence Results for Convex Maximization
$0\leq\langle\nabla f(x^{t}),x^{t+1}-x^{t}\rangle\leq f(x^{t+1})-f(x^{t})\leq f^{*}-f(x^{t})<+\infty,$
@@
$\langle\nabla f(x^{t}),x^{t+1}-x^{t}\rangle\geq 0\quad\forall t=0,1,\ldots,T.$
@@
$f(x^{t+1})\geq f(x^{t})+\langle\nabla f(x^{t}),x^{t+1}-x^{t}\rangle+\alpha D_{h}(x^{t+1}\,\|\,x^{t})\quad\forall\;t=0,1,\ldots,T.$
Summing Section 4.1 up over $t=0,\ldots,T$ and applying Section 4.1, we have
$\alpha T\min_{0\leq t\leq T}D_{h}(x^{t+1}\,\|\,x^{t})\leq\alpha\sum_{t=0}^{T}D_{h}(x^{t+1}\,\|\,x^{t})\leq f(x^{T+1})-f(x^{0})\leq f^{*}-f(x^{0}).$
@@
$D_{IS}(y\,\|\,x)=D_{h}(y\,\|\,x)=\sum_{i}\left(-\log{\frac{y_{i}}{x_{i}}}+\frac{y_{i}}{x_{i}}-1\right),$ (11)
where $D_{IS}(\cdot\|\,\cdot)$ refers to the Itakura–Saito distance. Then, for the $\log$-type function in Eq. 10, we trivially have
$g(y)\geq g(x)+\langle\nabla g(x),y-x\rangle+\min_{i}a_{i}\;D_{IS}(y\|x),$ (12)
@@
$\frac{1}{3}\sum_{i}\left(\frac{y_{i}}{x_{i}}-1\right)^{2}\leq D_{IS}(y\,\|\,x)$ (13)
if $-\log{\frac{y_{i}}{x_{i}}}+\frac{y_{i}}{x_{i}}-1\leq c:=\sqrt{3}-1-\frac{1}{2}\log{3}$ for all $i$.
@@
4.2 Convergence to Approximate Competitive Equilibrium
@@
We prove this by contradiction. Assume there exists a $j\in[m]$, such that $p^{t}_{j}=0$. Then note that $p^{t}_{j}/d_{ij}<\beta^{t}_{i}$ for all $i\in[n]$ (this follows from the fact that every $\beta^{t}_{i}$ is strictly positive). Due to the strict inequality, there must exist some (possibly small) $\gamma>0$ such that if we increase $p_{j}^{t}$ by $\gamma$, and decrease the price of every other good by $\gamma/(\sum_{i}B_{i}-1)$, then we still have a feasible set of prices $p^{\prime}$, and $\max_{j\in[m]}p^{t}_{j}/d_{ij}$ strictly decreases for all $i\in[n]$. Let $\beta^{\prime}_{i}=\max_{j\in[m]}p_{j}^{\prime}/d_{ij}$. Note that $\beta^{\prime}_{i}<\beta^{t}_{i}$ for all $i\in[n]$ and thus $\sum_{i\in[n]}B_{i}\cdot\beta^{\prime}_{i}/\beta^{t-1}_{i}<\sum_{i\in[n]}B_{i}\cdot\beta^{t}_{i}/\beta^{t-1}_{i}$ which contradicts the optimality of $\beta^{t}$ when solving the GFW LP. ∎
Lemma 2.
@@
$\min_{p,\beta\geq 0}$ $\sum_{i}B_{i}\frac{\beta_{i}}{\beta^{t-1}_{i}}$ (14)
@@
Let $x$ and $\mu$ denote the dual variables corresponding to the first and second constraints of Eq. 14, respectively. We use $(p^{t},\beta^{t})$ and $(x^{t},\mu^{t})$ to denote the optimal primal and dual variables at the $t$-th iteration, respectively.
By KKT conditions and strong duality of LP, $(p,\beta)$ is an optimal solution to Eq. 14 if and only if there exists a pair of $(x,\mu)$ such that $(p,\beta,x,\mu)$ satisfies
@@
Recall that $(p^{t},\beta^{t})$ is an optimal solution to Eq. 14. By condition (ii), we have $\frac{p^{t}_{j}}{d_{ij}}=\beta^{t}_{i}\geq\frac{p^{t}_{j^{\prime}}}{d_{ij^{\prime}}}$ for all $j^{\prime}\in[m]$ if $x^{t}_{ij}>0$ for all $i,j$. This guarantees every buyer is allocated her optimal bundle subject to the amount of budget she earns.
@@
$\sum_{i}x^{t}_{ij}=-\mu=\frac{\sum_{i}B_{i}\frac{\beta_{i}}{\beta^{t-1}_{i}}}{\sum_{i}B_{i}}\quad\text{if }p^{t}_{j}>0.$ (17)
By 1, this holds for all $j\in[m]$. It follows that
$\sum_{i}\bar{x}^{t}_{i}=\frac{\sum_{i}B_{i}}{\sum_{i}B_{i}\frac{\beta^{t}_{i}}{\beta^{t-1}_{i}}}\sum_{i}x^{t}_{i}=1\quad\mbox{for all }j\in[m].$ (18)
@@
Therefore, the lemma follows by the definition of $\epsilon$-strongly approximate CE. ∎
@@
Therefore, by Lemmas 2 and 6, we can conclude that the GFW algorithm finds an $\epsilon$-approximate CEEI in $\left\lceil\frac{3(g^{*}-g(x^{0}))}{\min_{i}a_{i}}\frac{1}{\epsilon^{2}}+\frac{g^{*}-g(x^{0})}{c\min_{i}a_{i}}\right\rceil$ iterations. Observe that $g^{*}-g(x^{0})$ can be upper bounded by
$\sup_{p,\beta\in Y}-\sum_{i}B_{i}\log(\beta_{i}).$
To this end, note that for each agent $i$, $1/\beta_{i}\leq\min_{j}d_{ij}/p_{j}$ (implied by the constraints of our program). Since $\sum_{j}p_{j}=\sum_{i}B_{i}$, this implies that there exists at least one chore $j$, such that $p_{j}\geq\sum_{i}B_{i}/m$, implying that $d_{ij}/p_{j}\leq mD/(\sum_{i}(B_{i}))$, where $D=\max_{i,j}d_{ij}$. Therefore, we have
$-\sum_{i}B_{i}\log(\beta_{i})=\sum_{i}B_{i}\log(1/\beta_{i})\leq\sum_{i}B_{i}\log(mD/(\sum_{i}B_{i}))\leq n\max_{i}B_{i}\log(mD/\min_{i}B_{i}).$
Theorem 7 (Complexity of finding an approximate CE).
The GFW algorithm finds an $\epsilon$-strongly approximate CE in $\mathcal{O}(n\frac{\max_{i}B_{i}}{\min_{i}B_{i}}\log(\frac{mD}{\min_{i}B_{i}})/\epsilon^{2})$ iterations.
@@
For CEEI, our results (Theorem 7) improve the current state-of-the-art convergence rate in terms of the number of iterations relative to $\epsilon$: The EPM by Boodaghians et al. (2022) requires $\mathcal{\tilde{O}}(n^{3}/\epsilon^{2})$ iterations, and the combinatorial algorithm by Chaudhury et al. (2022) requires $\mathcal{\tilde{O}}(nm/\epsilon^{2})$ iterations, in contrast to $\mathcal{{\tilde{O}}}(n/\epsilon^{2})$ iterations of GFW [^5] For CEEI, note that $\max_{i}B_{i}=\min_{i}B_{i}=1$., to reach an approximate CE. Furthermore, we expect the cost of each iteration to be cheaper for GFW. In GFW, we only need to solve an LP in every iteration. Moreover, the approximate solution of the LP poses no numerical issues. In contrast, solving the convex QP in EPM is a more challenging task. It is hard to find exact running times for general convex QPs, and even quite good approximate solutions can lead to serious numerical issues, as we show in Section 5. Further, we believe that the LP solved in every iteration of GFW can be solved (in worst case) using algorithms faster than the ones used for general-purpose LPs van den Brand (2020), and we leave this as an interesting question for future research.
@@
The result in Theorem 7 can also be recovered with mildly stronger guarantees through a primal-dual analysis of the GFW (See Remark Remark). The analysis in Remark Remark gives us further insights on each run of GFW: for instance, in each iteration of GFW where we do not yet have a good approximate CE point, we make a large improvement on the objective function (5).
5 Experiments